How do I solve this:
200/r+10=(200/r) - 1
r is supposed to =40
Thank you!

Answer by stanbon(75887)

How do I solve this:
200/(r+10) = (200/r) - 1
---
Multiply thru by r(r+10) to get:
200r = 200(r+10) -r(r+10)
---
200r = 200r + 2000 - r^2 - 10r
r^2 + 10r + 2000 = 0
Factor:
(r+50)(r-40) = 0
---
r = -50 or r = 40
========================
Cheers,
Stan H.
================